# Detect, Explain, Escalate: Sustainable Dialogue Breakdown Management for LLM Agents

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated substantial capabilities in conversational AI applications, yet their susceptibility to dialogue breakdowns poses significant challenges to deployment reliability and user trust. This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing resource-efficient operation. Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown detector and explainer. This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes to the BETOLD dataset, improving accuracy by 7% over its baseline. (2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment. These are integrated into an “escalation” architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and computational overhead. Our fine-tuned model and prompting strategies achieve state-of-the-art performance on DBDC5 and strong results on BETOLD, outperforming specialized classifiers on DBDC5 and narrowing the performance gap to larger proprietary models. The proposed monitor–escalate pipeline reduces inference costs by 54%, providing a cost-effective and interpretable solution for robust conversational AI in high-impact domains. Code and models will be publicly released.
@@ I Introduction
Conversational artificial intelligence (AI) has experienced substantial growth driven by advances in large language models (LLMs) such as ChatGPT [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama]. ChatGPT’s rapid adoption, reaching 100 million users within two months [Yatoo2023ChatGPT, Khosravi2023Chatbots], demonstrates the widespread interest in LLM-based conversational systems and their potential to transform human-computer interaction. These models are now embedded in applications spanning customer service, content creation, marketing, and education, enabling new forms of human–AI interaction [Currie2023A, Gabriella2024The, Arviani2023ChatGPT].
@@
Despite these advancements, the accelerated integration of LLMs into critical domains has revealed significant challenges in dialogue breakdown detection and mitigation, which directly impact user trust and conversational effectiveness [Ghassel2024Are, Fan2022Building, candello_recovering_2018]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [higashinaka_dialogue_2016, Ashktorab]. Table I illustrates a common breakdown pattern, where the assistant’s utterances contradict one another.
As LLM-based systems are deployed in high-stakes environments [Currie2023A, Gabriella2024The, Arviani2023ChatGPT], robust handling of such breakdowns becomes increasingly important. The tendency of LLMs to produce overly confident yet inaccurate or hallucinated outputs further complicates safe deployment [xiong_can_2023, zhou_larger_2024]. There is therefore a pressing need for methods that not only detect dialogue breakdowns reliably but also support mitigation strategies that preserve conversational quality and user trust.
Traditionally, dialogue breakdown detection has been approached using specialized classifiers trained on labelled datasets such as the Dialogue Breakdown Detection Challenge (DBDC) [higashinaka_dialogue_2016, tsunomori-etal-2018-evaluating]. While these models achieve strong benchmark performance, their generalization to diverse, real-world contexts remains limited. At the same time, recent work shows that general-purpose LLMs still lag behind human performance for nuanced conversational behavior, despite their broad coverage and world knowledge [finch_leveraging_2023, finch_dont_2023]. This gap indicates that neither specialized classifiers nor generalist LLMs alone provide a complete solution.
@@
First, we fine-tune a parameter-efficient Llama-3.1 8B model [llama] on both English and Japanese tracks of the DBDC5 dataset [Higashinaka2020DBDC5], augmenting supervision with synthetic reasoning trajectories generated by a larger teacher model, Llama-3.3 70B. These distilled reasoning traces are used during training to shape the student model’s decision process and to support interpretable justifications at inference time. We then evaluate the model’s generalization on the BETOLD dataset [terragni2022_BETOLD], which focuses on task-oriented dialogue breakdowns in realistic customer-service settings.
Second, we conduct a comprehensive comparative analysis of both closed-source frontier models (OpenAI [openai_gpt-4_2023], Anthropic [ClaudeModelCard]) and open-source alternatives (Meta [llama], Mistral [jiang_mixtral_2024], DeepSeek [deepseek]) on dialogue breakdown detection. We explore and propose novel prompting schemes, including few-shot learning [brown_language_2020, palm2023], chain-of-thought prompting [wei_chain--thought_2023], and analogical reasoning with curriculum learning [analogical, xu-etal-2020-curriculum, Ryu2024CurricuLLM:], to elicit more systematic reasoning while regulating token costs.
@@ II Related Work
Ensuring robustness in conversational AI systems, particularly for dialogue breakdown detection and mitigation, is a central research challenge. Dialogue breakdowns occur when coherence or relevance is disrupted, hindering the smooth progression of a conversation and reducing user satisfaction [higashinaka_dialogue_2016, Ashktorab]. Practical conversational agents must therefore both detect breakdowns and employ recovery strategies to maintain engagement and trust [mctear, skantze-2017-towards].
@@ II-A Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has produced specialized classifiers that perform strongly on benchmarks such as DBDC5 (see Section V-A[higashinaka_dialogue_2016]. Leading approaches rely on fine-tuned transformer encoders [attentionIsAll]. For example, the best-performing model on the DBDC5 English track, BERT+SSMBA [Devlin2019BERT, ng_improving_2020], augments a BERT-based classifier with unlabelled dialogue data through extended pre-training on dialogue-rich corpora (e.g., Reddit) and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [ng_ssmba_2020]. This design improves robustness by exploiting large quantities of unlabelled conversational data.
Semi-supervised approaches further advance the state of the art. The S2T2 model employs a dual-teacher paradigm that leverages both labelled and unlabelled dialogues [lin_semi-supervised_2022]. One teacher is trained on high-quality labelled data, while the other is trained on masked dialogue variants. A student model is jointly guided by both teachers. S2T2 reaches new state-of-the-art performance on DBDC5 using RoBERTa-large on the English track and XLM-R-large with a context-matching mechanism on the Japanese track [liu2019roberta, conneau2020unsupervised, lin_semi-supervised_2022].
More recently, general-purpose LLMs have been evaluated for dialogue breakdown detection. Finch et al. [finch_leveraging_2023, finch_dont_2023] studied ChatGPT’s behaviour across nine categories in the ABC-Eval dataset. While ChatGPT outperforms specialized models on some categories (e.g., empathetic behaviour), it still falls short of human performance on many dialogue tasks. This suggests that current large models, despite their sophisticated capabilities, remain unreliable for fine-grained breakdown detection. To our knowledge, prior work has not systematically evaluated multiple LLM families on dialogue breakdown detection and remediation.
@@ II-B Current State of Conversational Agents
Beyond detection, modern conversational agents increasingly incorporate mechanisms to mitigate or recover from breakdowns. In the DBDC5 recovery track, for instance, systems respond to detected breakdowns by asking clarifying questions or providing corrections [higashinaka_dialogue_2016]. Industrial systems such as ChatGPT and Claude are trained with alignment techniques like reinforcement learning from human feedback (RLHF) to reduce toxic, incoherent, or nonsensical outputs [openai_gpt-4_2023, ClaudeModelCard]. Claude-2, for example, uses internal debate and self-critique during training to identify and minimize reasoning flaws [claude2]. These mechanisms reduce the incidence of breakdowns but do not eliminate them, especially under distribution shift or adversarial inputs.
@@ II-C Techniques in Conversational AI
Analogical Reasoning. Analogical prompting encourages a model to draw on relevant past examples by analogy when solving a new problem. Instead of relying on hand-crafted exemplars, the model is prompted to generate analogous examples in context and then solve the target query [analogical, webb_emergent_2023]. This approach can improve reasoning by creating problem-specific exemplars without additional labelling.
Chain-of-Thought Reasoning. Chain-of-thought (CoT) prompting instructs the model to produce intermediate reasoning steps before giving a final answer [zs-cot, fs-cot, wei_chain--thought_2023]. CoT has been shown to improve performance on arithmetic, logical, and commonsense reasoning tasks. For dialogue agents, CoT can be applied internally to reason about user intent, dialogue history, and knowledge bases, thereby reducing non sequiturs and hallucinations.
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLMs excel at in-context learning, adapting to new tasks given instructions and a small number of examples. In few-shot learning, a prompt includes a handful of example dialogues or question–answer pairs, and the model generalizes the pattern to new inputs. GPT-3 demonstrated that large models can function as few-shot learners without task-specific fine-tuning [brown_language_2020]. For dialogue tasks, this allows rapid adaptation to new domains using carefully crafted prompts rather than retraining.
@@ II-D Hierarchical Architectures
Hierarchical or multi-tiered architectures are a common strategy for balancing performance and cost in AI systems [chen2024frugalgpt, yue2024large]. In the context of LLMs, such architectures typically use smaller, faster models for routine inference and reserve larger, more capable models for complex or high-stakes cases [10.5555/3692070.3693614, 10.5555/3586589.3586709]. This can be viewed as a system-level “mixture of experts” [2017olln], where routing decisions balance accuracy and computational expense. Our work extends this paradigm for real-time dialogue management, positioning a lightweight monitor in front of more capable models to create a practical and sustainable framework for conversational agents.
@@ III Problem Definition
Dialogue breakdown refers to the deterioration of coherence, relevance, or conversational fluency between a user and a conversational agent [higashinaka_dialogue_2016, skantze-2017-towards]. Breakdowns may appear as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that disrupt the dialogue and erode user trust [Fan2022Building, candello_recovering_2018]. As LLMs from the OpenAI [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama] families are deployed in diverse conversational tasks, detecting and mitigating such breakdowns becomes critical. This is exacerbated by LLMs’ propensity to produce confident but incorrect or hallucinated content [xiong_can_2023, zhou_larger_2024].
@@ III-B Consolidation of Three-Class Annotations
Following prior works [Ghassel2024Are, lin_semi-supervised_2022], we preprocess DBDC5’s three-way labels (Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB)) into a binary Breakdown/Non-Breakdown scheme. Ambiguous annotations (PB) are merged into the Breakdown class. When multiple annotators label each system turn, let $p_{i}$ be the fraction who assign B or PB to utterance $s_{i}$. We define a binary label $b_{i}$ by thresholding:
@@ III-C Conversation-Level Labelling
In task-oriented datasets such as BETOLD [terragni2022_BETOLD], labels are assigned at the conversation level. Each dialogue $D$ is marked as a failure if, for example, the user hangs up or requests escalation to a human agent. We denote this label by:
@@ IV-A Supervised Fine-tuning with Reasoning Augmentation
To support real-time monitoring, we fine-tune a smaller model on labelled breakdown data using supervised fine-tuning (SFT). We select Llama-3.1 8B [llama] as a balance between accuracy and computational cost. The model is fine-tuned on the DBDC5 English and Japanese tracks [higashinaka_dialogue_2016] for per-utterance breakdown labels.
@@
We augment the training inputs with these synthetic reasoning traces, forming an enriched dataset $\mathcal{D}^{\prime}$. The student model $S$ is fine-tuned on $\mathcal{D}^{\prime}$ to (i) predict the binary breakdown label $b_{i}$ by minimizing cross-entropy loss and (ii) generate a textual justification $\hat{j}_{i}$. Teacher-generated traces $r_{i}$ serve as targets for justification, encouraging the student to internalize structured reasoning patterns that improve both classification and interpretability. The cross-entropy loss [shannon_cel] is:
@@ IV-B Prompting Strategies
Chain-of-Thought (CoT). CoT prompts encourage multi-step reasoning by instructing $\mathcal{G}$ to “think step by step” before deciding:
$\Pi_{\mathrm{CoT}}(\mathcal{H}_{i},s_{i})\;=\;\bigl[\mathcal{H}_{i},\,s_{i},\,\text{``Let’s think step by step.''}\bigr].$
@@
$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl\{\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr\}\;\cup\;(\mathcal{H}_{i},s_{i}).$
@@
$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}.$
@@ IV-C Deployment Architecture
    AI Assistant: A base assistant model generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and history $\mathcal{H}_{i}$:
@@
    $s_{i}^{*}:=\mathcal{G}_{\mathrm{superior}}\!\bigl(\mathcal{H}_{i},u_{i},\hat{j}_{i}\bigr),$

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated substantial capabilities in conversational AI applications, yet their susceptibility to dialogue breakdowns poses significant challenges to deployment reliability and user trust. This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing resource-efficient operation. Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown detector and explainer. This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes to the BETOLD dataset, improving accuracy by 7% over its baseline. (2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment. These are integrated into an “escalation” architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and computational overhead. Our fine-tuned model and prompting strategies achieve state-of-the-art performance on DBDC5 and strong results on BETOLD, outperforming specialized classifiers on DBDC5 and narrowing the performance gap to larger proprietary models. The proposed monitor–escalate pipeline reduces inference costs by 54%, providing a cost-effective and interpretable solution for robust conversational AI in high-impact domains. Code and models are publicly available. [^1] https://github.com/aghassel/dee-dialogue-agents
@@ I Introduction
Conversational artificial intelligence (AI) has experienced substantial growth driven by advances in large language models (LLMs) such as ChatGPT [38], Claude [2], and Llama [44]. ChatGPT’s rapid adoption, reaching 100 million users within two months [53, 29], demonstrates the widespread interest in LLM-based conversational systems and their potential to transform human-computer interaction. These models are now embedded in applications spanning customer service, content creation, marketing, and education, enabling new forms of human–AI interaction [11, 20, 4].
@@
Despite these advancements, the accelerated integration of LLMs into critical domains has revealed significant challenges in dialogue breakdown detection and mitigation, which directly impact user trust and conversational effectiveness [21, 15, 7]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [23, 5]. Table I illustrates a common breakdown pattern, where the assistant’s utterances contradict one another.
As LLM-based systems are deployed in high-stakes environments [11, 20, 4], robust handling of such breakdowns becomes increasingly important. The tendency of LLMs to produce overly confident yet inaccurate or hallucinated outputs further complicates safe deployment [50, 55]. There is therefore a pressing need for methods that not only detect dialogue breakdowns reliably but also support mitigation strategies that preserve conversational quality and user trust.
Traditionally, dialogue breakdown detection has been approached using specialized classifiers trained on labelled datasets such as the Dialogue Breakdown Detection Challenge (DBDC) [23, 45]. While these models achieve strong benchmark performance, their generalization to diverse, real-world contexts remains limited. At the same time, recent work shows that general-purpose LLMs still lag behind human performance for nuanced conversational behavior, despite their broad coverage and world knowledge [18, 17]. This gap indicates that neither specialized classifiers nor generalist LLMs alone provide a complete solution.
@@
First, we fine-tune a parameter-efficient Llama-3.1 8B model [44] on both English and Japanese tracks of the DBDC5 dataset [24], augmenting supervision with synthetic reasoning trajectories generated by a larger teacher model, Llama-3.3 70B. These distilled reasoning traces are used during training to shape the student model’s decision process and to support interpretable justifications at inference time. We then evaluate the model’s generalization on the BETOLD dataset [43], which focuses on task-oriented dialogue breakdowns in realistic customer-service settings.
Second, we conduct a comparative analysis of both closed-source frontier models (OpenAI [38], Anthropic [2]) and open-source alternatives (Meta [44], Mistral [28], DeepSeek [12]) on dialogue breakdown detection. We explore and propose novel prompting schemes, including few-shot learning [6, 9], chain-of-thought prompting [49], and analogical reasoning with curriculum learning [52, 51, 39], to elicit more systematic reasoning while regulating token costs.
@@ II Related Work
Ensuring robustness in conversational AI systems, particularly for dialogue breakdown detection and mitigation, is a central research challenge. Dialogue breakdowns occur when coherence or relevance is disrupted, hindering the smooth progression of a conversation and reducing user satisfaction [23, 5]. Practical conversational agents must therefore both detect breakdowns and employ recovery strategies to maintain engagement and trust [34, 42].
@@ II-A Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has produced specialized classifiers that perform strongly on benchmarks such as DBDC5 (see Section V-A[23]. Leading approaches rely on fine-tuned transformer encoders [46]. For example, the best-performing model on the DBDC5 English track, BERT+SSMBA [13, 36], augments a BERT-based classifier with unlabelled dialogue data through extended pre-training on dialogue-rich corpora (e.g., Reddit) and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [35]. This design improves robustness by exploiting large quantities of unlabelled conversational data.
Semi-supervised approaches further advance the state of the art. The S2T2 model employs a dual-teacher paradigm that leverages both labelled and unlabelled dialogues [31]. One teacher is trained on high-quality labelled data, while the other is trained on masked dialogue variants. A student model is jointly guided by both teachers. S2T2 reaches new state-of-the-art performance on DBDC5 using RoBERTa-large on the English track and XLM-R-large with a context-matching mechanism on the Japanese track [32, 10, 31].
More recently, general-purpose LLMs have been evaluated for dialogue breakdown detection. Finch et al. [18, 17] studied ChatGPT’s behaviour across nine categories in the ABC-Eval dataset. While ChatGPT outperforms specialized models on some categories (e.g., empathetic behaviour), it still falls short of human performance on many dialogue tasks. This suggests that current large models, despite their sophisticated capabilities, remain unreliable for fine-grained breakdown detection. To our knowledge, prior work has not systematically evaluated multiple LLM families on dialogue breakdown detection and remediation.
@@ II-B Current State of Conversational Agents
Beyond detection, modern conversational agents increasingly incorporate mechanisms to mitigate or recover from breakdowns. In the DBDC5 recovery track, for instance, systems respond to detected breakdowns by asking clarifying questions or providing corrections [23]. Industrial systems such as ChatGPT and Claude are trained with alignment techniques like reinforcement learning from human feedback (RLHF) to reduce toxic, incoherent, or nonsensical outputs [38, 2]. Claude-2, for example, uses internal debate and self-critique during training to identify and minimize reasoning flaws [3]. These mechanisms reduce the incidence of breakdowns but do not eliminate them, especially under distribution shift or adversarial inputs.
@@ II-C Techniques in Conversational AI
Analogical Reasoning. Analogical prompting encourages a model to draw on relevant past examples by analogy when solving a new problem. Instead of relying on hand-crafted exemplars, the model is prompted to generate analogous examples in context and then solve the target query [52, 48]. This approach can improve reasoning by creating problem-specific exemplars without additional labelling.
Chain-of-Thought Reasoning. Chain-of-thought (CoT) prompting instructs the model to produce intermediate reasoning steps before giving a final answer [30, 47, 49]. CoT has been shown to improve performance on arithmetic, logical, and commonsense reasoning tasks. For dialogue agents, CoT can be applied internally to reason about user intent, dialogue history, and knowledge bases, thereby reducing non sequiturs and hallucinations.
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLMs excel at in-context learning, adapting to new tasks given instructions and a small number of examples. In few-shot learning, a prompt may include example dialogues or question–answer pairs, and the model generalizes the pattern to new inputs. GPT-3 demonstrated that large models can function as few-shot learners without task-specific fine-tuning [6].
@@ II-D Hierarchical Architectures
Hierarchical or multi-tiered architectures are a common strategy for balancing performance and cost in AI systems [8, 54]. In the context of LLMs, such architectures typically use smaller, faster models for routine inference and reserve larger, more capable models for complex or high-stakes cases [37, 16]. This can be viewed as a system-level “mixture of experts” [41], where routing decisions balance accuracy and computational expense. Our work extends this paradigm for real-time dialogue management, positioning a lightweight monitor in front of more capable models to create a practical and sustainable framework for conversational agents.
@@ III Problem Definition
Dialogue breakdown refers to the deterioration of coherence, relevance, or conversational fluency between a user and a conversational agent [23, 42]. Breakdowns may appear as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that disrupt the dialogue and erode user trust [15, 7]. As LLMs from the OpenAI [38], Claude [2], and Llama [44] families are deployed in diverse conversational tasks, detecting and mitigating such breakdowns becomes critical. This is exacerbated by LLMs’ propensity to produce confident but incorrect or hallucinated content [50, 55].
@@ III-B Consolidation of Three-Class Annotations
Following prior works [21, 31], we preprocess DBDC5’s three-way labels (Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB)) into a binary Breakdown/Non-Breakdown scheme. Ambiguous annotations (PB) are merged into the Breakdown class. This specific binarization strategy aligns with the evaluation protocol employed by the S2T2 baseline [31], ensuring that the performance comparisons reported in Section VI are methodologically valid.
When multiple annotators label each system turn, let $p_{i}$ be the fraction who assign B or PB to utterance $s_{i}$. We define a binary label $b_{i}$ by thresholding:
@@ III-C Conversation-Level Labelling
In task-oriented datasets such as BETOLD [43], labels are assigned at the conversation level. Each dialogue $D$ is marked as a failure if, for example, the user hangs up or requests escalation to a human agent. We denote this label by:
@@ IV-A Supervised Fine-tuning with Reasoning Augmentation
To support real-time monitoring, we fine-tune a smaller model on labelled breakdown data using supervised fine-tuning (SFT). We select Llama-3.1 8B [44] as a balance between accuracy and computational cost. The model is fine-tuned on the DBDC5 English and Japanese tracks [23] for per-utterance breakdown labels.
@@
We augment the training inputs with these synthetic reasoning traces, forming an enriched dataset $\mathcal{D}^{\prime}$. The student model $S$ is fine-tuned on $\mathcal{D}^{\prime}$ to (i) predict the binary breakdown label $b_{i}$ by minimizing cross-entropy loss and (ii) generate a textual justification $\hat{j}_{i}$. Teacher-generated traces $r_{i}$ serve as targets for justification, encouraging the student to internalize structured reasoning patterns that improve both classification and interpretability. The cross-entropy loss [40] is:
@@ IV-B Prompting Strategies
Chain-of-Thought (CoT). CoT encourages multi-step reasoning by instructing $\mathcal{G}$ to “think step by step” before deciding:
$\Pi_{\mathrm{CoT}}(\mathcal{H}_{i},s_{i})\;=\;\bigl(\mathcal{H}_{i},\,s_{i},\,\text{``Let’s think step by step.''}\bigr).$
@@
$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl(\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime}),\,(\mathcal{H}_{i},s_{i})\bigr).$
@@
$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\bigl(\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\bigr).$
@@ IV-C Deployment Architecture
    AI Assistant: An assistant model generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and history $\mathcal{H}_{i}$:
@@
    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl(\mathcal{H}_{i},u_{i},\hat{j}_{i}\bigr),$